OpenStudy (jessiegonzales):

MEDAL!!! Solve the inequality. Graph the solution. 8. lx+3/4l <2

OpenStudy (jessiegonzales):

A. x<-1 or x>-1/2 B. -5/32<x<5/32 C. -1<X<-1/2 D. X<-5/32 OR X>5/32

OpenStudy (jessiegonzales):

@nincompoop

OpenStudy (anonymous):

it s C

OpenStudy (anonymous):

if you want details just tell me and i ll send

OpenStudy (jessiegonzales):

DETAILS PLEASE

OpenStudy (anonymous):

\[\left| x+\frac{ 3 }{ 4 } \right|<2\] means (x + 3/4) is between -2 and 2, so you need to solve this compound inequality. \[-2<x+\frac{ 3 }{ 4}<2\]

OpenStudy (anonymous):

Do you know how to get x by itself in the middle?

OpenStudy (jessiegonzales):

divide?

OpenStudy (anonymous):

3/4 is added to x, so you need to subtract